As Gemini South resumed nighttime operation, GMOS-S was checked out during last weekend and started operating in queue mode. To be noted is that the CCD1 CTE problem reported on Jan10, 2020 is still present, meaning that programs using the N&S mode for spectroscopy are potentialliy affected, as well as IFU programs. These programs will remain on hold, with the exception of programs where this issue has no impact on the science (e.g. region of interest outside of CCD1)
After the COVID-19 telescope shutdown, F2 will continue off the telescope until further notice, due to pending work on the instrument cold heads.
GPI is not available in semester 20B or later due to being removed from science operations for the GPI 2.0 upgrade and move to Gemini North.
GMOS-N data taken between the nights of 31 August 2020 and 8 September 2020 show a new flat field feature from dust in the optics, as marked in the image below. A preliminary cleaning of the accessible GMOS optics performed during daytime on 3 September 2020 was not successful in removing the new dust feature and the overall effects were found to be negligible.
GPI is now warming up since Monday the 16th due to the unplanned shutdown due to the Covid-19 outbreak. The instrument is now safe in a shutdown state on the telescope.
As Gemini South is closed for operations from last Monday due to the Covid-19 outbreak (see https://www.gemini.edu/news/gemini-south-shutdown-until-further-notice), GMOS-S has been warmed up and switched off. This shutdown will last until further notice
Zorro has now an H-alpha filter! Addittionally, the same new control software has also been implemented. The first Zorro run of the semester is ongoing.
F2 will be off the telescope until March 27th, due to work on the instrument cold heads.
GPI is not offered in Semester 2020B as the instrument will leave GS in early 20B for upgrades and later commissioning at GN.
The Gemini North Adaptive Optics system, Altair, is currently offline due to problems with its deformable mirror (DM) electronics. Issues with other mechanisms on the optical bench have been resolved, however, we are still working on finding the root cause of the DM electronics failure. We are making progress but still need to test a number of components. We will have more updates by the end of February. Please email your contact scientist if you have any questions or concerns about the potential impact on your program.